# Design of superparamagnetic nanoparticle-materials for high-frequency inductor cores

### v3

@@ Abstract
The progress in the semiconductor industry has resulted in great demand for high-frequency magnetic materials applicable in microfabricated inductor cores. Nanocomposite materials, containing magnetic nanoparticles in a non-conducting matrix, may provide a solution for materials with high susceptibility (/permeability) and low power loss in the MHz regime, where traditional ferrites fail in performance. Here, we present a design guide for usage of magnetic nanoparticles in such materials. We use statistical mechanics methods to derive the magnetic susceptibility of nanoparticles in case of uniaxial or cubic anisotropy, as function of particle size and applied field direction, and investigate shape and interaction effects on the susceptibility. Using the derived susceptibilities, with inductor-core applications in mind, we show that close-to-spherical particles of materials with high saturation magnetization and low magnetic anisotropy, such as FeNi3, are optimal. Additionally, the particle size shall be optimized to be as large as possible while maintaining superparamagnetic behaviour at the relevant frequency. Based on this, we predict that high particle susceptibilities of 700 (/1500) are possible for randomly oriented (/aligned) 201 nm diameter FeNi3 particles, together with high-frequency stability, shown by low out-of-phase component at 2 MHz. This implies that materials containing nanoparticles have the potential to be tuned to outperform state-of-the-art ferrite inductor-core materials at MHz-frequencies.
@@ I Introduction
Magnetic components, such as inductors and transformers, are essential for power electronics in many handheld devices. However, realisation of efficient micro-inductors is challenging as inductance scales inversely with size [1, 2]. Although a decrease in inductance can be counteracted by higher operating frequencies, magnetic materials available today become inefficient and heat up rapidly at elevated frequencies due to eddy-current losses [3, 4, 5]. Miniaturisation of magnetic components in electronics is therefore limited by the performance of soft magnetic materials [6]. Several roadmaps for power electronics, identifies the lack of suitable magnetic materials as (one of) the major obstacle(s) for achieving smaller, faster, greener, and more efficient electronic devices [1, 6].
@@
The challenge lies in achieving sufficiently high susceptibility while avoiding significant power losses from eddy currents and magnetic hysteresis at high operation frequencies. To this end, for a magnetic material to be used as a core for micro-inductors, its in-phase susceptibility, , needs to be above 50-100 (the turqouise region shown in Fig. 1), while losses should remain below 200 mW/cm3 at an operating frequency of 2 MHz and a flux density of 30 mT [3]. If losses are reduced below 20 mW/cm3, a -range of 20-50 would be acceptable [1, 6].
@@
For today’s power converters, operating at up to about 0.5 MHz, the typical core materials are sintered, fine-grained ceramics of ferrites (MnZnNiFe-oxides), with poor conductivity, which limits eddy current losses at lower frequencies. The properties of the ferrites vary with their grain sizes and material composition. For instance, the commonly used ferrite, TDK PC200, has high susceptibility of 800 (shown in Fig. 1) but only up 1-2 MHz, where eddy current losses eventually increases drastically, and the susceptibility drops [3]. New ferrite materials, like Fair-Rite 67 (also shown in Fig. 1), allow for higher frequency operation, but have lower susceptibility [4]. However, so far, attempts to make inductor core materials to cover applications in the high-frequency-high-susceptibility regime (the turquoise area of interest in Fig. 1) have not been successful [5] and therefore the field of power electronics lacks alternative material solutions.
@@
Magnetic composites using nanoparticles have recently emerged as research materials [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], but their potential for inductor materials has not yet well been fully considered. Notable reasons for using nanoparticles are: 1) the magnetic particles, even if metallic, are so small that eddy current are neglible, and can be electrically insulated by a non-conductive matrix, 2) the possibly high susceptibility and low hysteresis within their superparamagnetic regime could be exploited. Superparamagnetism arises when the magnetic moments of the nanoparticles fluctuate on faster timescales than the applied field. This depends on particle size, shape, anisotropy, temperature, applied field frequency, and also particle interactions [19].
@@
Several groups have reported susceptibility data for materials containing magnetic nanoparticles ( nm diameter), as presented in Fig. 1. The found susceptibilities are generally too low to match the area of interest for future power electronics (the turquoise area in Fig. 1). However, it is not clear if composites containing superparamagnetic particles could actually be able to reach the desired susceptibility in the MHz-range if they were optimized for it. Recently, it has been shown that the susceptibility of single-domain nanoparticles is not limited by their shape (in contrast to soft multi-domain micro-particles) [Zambach2025], hence there is in principle no upper limit for the susceptibility of such nanoparticles, but the actual susceptibility will depend on material parameters such as magnetic anisotropy and saturation magnetization as well as particle size.
@@
In order to evaluate the potential of nanoparticles for use in inductor core materials, we here present a framework for calculating the DC- and AC-susceptibility of single-domain magnetic nanoparticles, both superparamagnetic and blocked. We focus on the need for high particle susceptibility and derive the susceptibility, using reported material values for specific materials, and considering both uniaxial and cubic anisotropy, and the size, shape, and alignment of the nanoparticles. Our theoretical framework reproduces the results from the nanocomposites in Fig. 1, and based on this predictive power, it pushes forward a way to tune and optimize properties.
@@ II Theoretical Framework
We consider a stationary single-domain ferromagnetic particle embedded in a solid, non-magnetic matrix. For single-domain particles, atom spins align ferromagnetically and rotate coherently such that the magnetic moment of the particle is , where is the particle volume and its saturation magnetization.
@@ II.1 Energy considerations
We define energies used later for susceptibility derivations.
@@
For cubic anisotropy we orient the coordinate system based on the case where the cubic anisotropy constant is positive, K 0, i.e., 3 mutually orthogonal easy axes, see Fig. 2(b). One easy axis is set to the polar axis and the other two easy axis are along the direction where the azimuth is 0 and . The anisotropy energy can then be expressed as
@@
with an effective shape anisotropy constant, . and are the demagnetisation factors along the principal spheroid axes. is positive (/negative) for oblate (/prolate) spheroids, respectively, and is the angle between the magnetic moment and the longer (/shorter) principal spheroid axis. For most soft magnetic materials, shape anisotropy dominates over magneto-crystalline anisotropy if the length difference between the axes is larger than 5-10%.
@@
Here is the center-center distance of the two spheres, denotes it’s Cartesian components, and the subscript for the angles denote the particle number.
@@ II.2 Susceptibility calculations
We derive the susceptibility of single domain particles for the two limiting cases: blocked particles with no thermal agitation, and superparamagnetic particles, where the particles’ magnetic moments fluctuate rapidly. For sufficiently small single-domain particles, thermal energy can induce superparamagnetism, i.e. moment fluctuations. The characteristic timescale, , for relaxation between easy directions, depends on the anisotropy barrier over the thermal energy . Magnetic particles can thus be classified depending on the timescale for the superparamagnetic relaxation versus the experimental timescale.
For blocked particles, thermal fluctuations are slower than the experimental timescale. In case of uniaxial anisotropy, the susceptibility depends on the direction of the applied field with respect to the particle easy axis, cf. the Stoner-Wohlfarth model [20]. For randomly oriented particles, the orientation averaged particle susceptibility is
since and and the blocked susceptibilities are and [21]. Blocked particles with uniaxial anisotropy and easy axis aligned parallel to the applied field will have a square hysteresis loop with coercive field of H. For the perpendicularly aligned case, no hysteresis is observed. For blocked particles with cubic anisotropy, a hysteresis loop opening is always observed no matter the direction of the applied field wrt. The 3 or 4 easy axes and thus only a small low field susceptibility is expected [22, 23, 24].
@@
where the integral over indicates integration over all possible energy states . The relevant single particle energies presented in Eqs. (1)-(5) depend on moment direction, thus the integral in Eq. (8) should here be taken over all possible moment directions: . For shorter notation we define following energy ratios
@@
The mean magnetisation along the direction of the applied field is then . The susceptibility can be found as the derivative of the mean magnetisation wrt. the applied field amplitude. For the non-interacting, uniaxial anisotropy particle case, with the applied field at an angle to the anisotropy axis, one finds the superparamagnetic particle susceptibility to be
@@
Figure 3 shows a plot of Eq. (11) for different values of . From the figure it is seen that the susceptibility ranges from to 0 in case of large anisotropy (). For low anisotropy () the susceptibility of the uniaxial anisotropy particle goes towards the random case value for all applied field angles, .
The orientation average of eq. (11) is the well known paramagnetic susceptibility
For a non-interacting particle with cubic anisotropy we find no dependence of the susceptibility on easy axis direction wrt. the applied field (i.e. and ). The initial susceptibility for the cubic anisotropy case is always the same as for the randomly oriented uniaxial anisotropy case in the superparamagnetic regime (Eq. (13)), no matter the strength of the anisotropy axis. This can be explained by the symmetric distribution of energy minima along 3 (or 4) easy axes, which does not change the probability of the moment to point along any specific direction, contrary to the uniaxial case where there is 1 direction (or plane) that is statistically more likely.
Dipolar interactions can be included in the above derivations. For the two particle case, assuming low anisotropy particles (), we find that the susceptibility per particle in case of interactions, , when the particles are placed such that they form a chain, respectively, parallel or perpendicular, to the field direction, is
@@
The dependence of the susceptibility on applied field frequency and the crossover from superparamagnetic towards blocked regime can be expressed by the concept of AC-susceptibility [21, 25]. At small , where the anisotropy barrier is not changed by the applied field, the AC-susceptibility for an applied sinusoidal field with angular frequency , , can be related to the superparamagnetic particle susceptibility, , and the superparamagnetic relaxation time, . The AC susceptibility is thus found by a Debye model of the form
@@
For uniaxial anisotropy particles depends on the direction of the applied field wrt. the easy axis [27, 28, 21, 25]. Magnetic moment relaxation time perpendicular to the anisotropy axis, , is assumed to be short, on the timescale of the attempt time [19]. Relaxation time over the anisotropy barrier, , i.e. parallel to the anisotropy axis, is slower, and often described by Arrhenius-type expressions. The AC-susceptibility for randomly oriented uniaxial anisotropy particles, , can thus be expressed by the weighted sum of the AC-susceptibility parallel and perpendicular with the easy axis:
@@
From the out-of-phase component one can estimate the hysteresis loss per particle volume by for small applied field amplitudes . For low anisotropy particles the out-of-phase component of the AC-susceptibility is overestimated by the above and can be regarded as an upper bound [26].
@@
For cubic anisotropy, expressions similar to the parallel case of (17) exist, but with a lower effective anisotropy constant of for and for [29, 30, 31, 32].
@@
From the formulas derived/presented above, we will in the next section estimate the maximum susceptibilities of nanoparticles. We consider candidate materials to be used with applied field at frequencies of up to 10 MHz.
III Evaluation of particles for nanocomposites
@@
Figure 4 shows the effective in-phase susceptibility as function of particle diameter from the superparamagnetic regime to the blocked regime for randomly oriented, spherical nanoparticles. Particle materials used are maghemite (-Fe2O3), Fe, Ni, FeNi3, and FeCo. We have used a field frequency of 10 MHz and particle susceptibilities are calculated by use of Eqs. (11) and (16)-(17). We have used reported effective uniaxial anisotropy and saturation magnetisation values for nanoparticle materials [33, 34, 35, 36, 37], the used values for and are given in the figure caption. We have here ignored the dependence of saturation magnetisation and effective anisotropy constant on synthesis routine and particle size, and used values reported for 3-10 nm sized particles (21-50 nm for FeNi3). Also, effective uniaxial anisotropy constants reported for materials known to have cubic magneto-crystalline anisotropy are used.
@@
From Fig. 4 it is seen that particle susceptibility increases with particle size for small particles in the superparamagnetic regime. When the particles reach sizes where their superparamagnetic relaxation time is comparable to the timescale of the exciting field, the susceptibility shows a peak. For larger particles, the susceptibility declines towards the blocked value, which depends on particle saturation magnetisation and anisotropy, but not on the particle size, cf. Eq. (7). In Fig. 4, both FeCo and FeNi3 particles are seen to have effective particle susceptibilities above 100 in the superparamagnetic region.
@@
We find that nanoparticles with high saturation magnetization and low anisotropy are able to display the largest susceptibility, as their magnetic moments are larger and their transition to the blocked regime occurs at larger diameters. The superparamagnetic region towards the susceptibility peak is of particular interest for applications due to the potential combination of high-susceptibility and limited hysteresis loss. Hence, the optimal size for a given material is on the ascending part of the susceptibility curve, before reaching the maximum of the susceptibility. We suggest, based on Fig. 4, the use of ca. 9 nm FeCo particles or ca. 20 nm FeNi3 particles for high-susceptibility applications such as micro-inductor core materials.
@@
The susceptibility of around 130 for the blocked FeNi3 particles, as seen in Fig. 4, could in principle be of interest, as such susceptibility value may be sufficient for use in micro-inductors. However, use of blocked particles requires particle alignment as only the case will show no hysteresis losses.
@@
From Eqs. (16)-(17) it is seen that also the dynamic susceptibility depends on the direction of the particle wrt. the applied field. In Fig. 5 we show the frequency dependence of the in- and out-of-phase susceptibility for uniaxial anisotropy FeNi3 particles with log-normal distributed diameters of 201 nm for cases with anisotropy axes parallel, perpendicularly and randomly oriented wrt. the applied field. It is seen for the frequency range of 10 kHz to 10 MHz that the 20 nm FeNi particles aligned with their easy axis parallel to the applied field have an in-phase susceptibility of 1600, i.e., more than twice the value for the randomly oriented case (700) and more than six times the perpendicular ”hard-axis” case (230). At around 100 MHz the in-phase part drops and the out-of-phase component peaks for the parallel aligned and random cases. This is due to the fact that in this frequency range. For the random case the susceptibility drops towards the random orientation blocked susceptibility of Eq. (7). For the perpendicular case we observe a completely flat susceptibility up to the GHz regime where . For the perpendicular case we find that susceptibility does not depend on particle size, with values close to the blocked, aligned case of Eq. (7). Alignment of particles by easy/hard axis thus allows for tuning of the magnetic properties.
@@
Even though shape anisotropy can increase susceptibility for particles with their easy axis aligned to the applied field, cf. Eq. (11) and Fig. 3, the increase in anisotropy will increase relaxation time for such particles and thus the out-of-phase component (the power losses due to magnetic hysteresis) of the particle. This is illustrated in Fig. 6, where the normalised in- and out-of-phase components of the susceptibility are shown for 201 nm FeNi3 particles aligned with easy-axes along the applied field. We have assumed that shape and magneto-crystalline anisotropy are in the same direction, i.e. . From Fig. 6 it is seen that the increase in the in-phase component of the susceptibility due to increasing the shape anisotropy is small (0.3 % at ), while the increase in out-of-phase component is substantial (22 % at ). It is also observed that for larger shape anisotropy () the particle becomes blocked. Smaller particles, i.e. particles further below the blocking diameter at a given frequency, will show stronger increase in in-phase susceptibility, but the proportional large growth of the out-of-phase susceptibility will be the same as for the shown particles. From these results on FeNi3 particles, we conclude that it is best to use spherical particles, as shape anisotropy will increase the out-of-phase component of the susceptibility due to increased superparamagnetic relaxation time, resulting in larger losses per induced magnetisation as discussed in the next section.
@@
Hysteresis losses per particle can be calculated from the out-of-phase component of the susceptibility (Eqs. (11)-(17)). For nanocomposites containing superparamagnetic particles we assume a linear dependence of nanocomposite on the volume fraction and particle susceptibility. The nanocomposite susceptibility varies with volume fraction of particles, therefore the applied field strength, , needed to obtain a certain magnetic flux in the material varies too. Hence, for a given application with a required flux density, the loss might be larger for a low susceptibility material than for high susceptibility materials due to the need for larger applied field. Therefore, assuming no interaction and same particle characteristics, the losses will increase with decreasing volume fraction of particles when the nanocomposites generate equal magnetisation.

### v4

@@ Abstract
The progress in the semiconductor industry has resulted in great demand for high-frequency magnetic materials that can be applied to micro-fabricated inductor cores. Nanocomposite materials, containing magnetic nanoparticles in a non-conducting matrix, may provide a solution for materials with high susceptibility or permeability and low power loss in the MHz regime, where traditional ferrites fail in performance. Here, we present a design guide for usage of magnetic nanoparticles in such materials. With statistical mechanics methods, we derive the magnetic susceptibility of nanoparticles in case of uniaxial or cubic anisotropy, as a function of particle size and applied field direction. We also investigate the role of particle shape and interactions. Using the derived susceptibilities, with inductor core applications in mind, we show that near-spherical particles of materials with high saturation magnetization and low magnetic anisotropy, such as FeNi3, are optimal. Additionally, we find that the particle size shall be as large as possible while maintaining superparamagnetic behaviour at the relevant operation frequency. Based on this, we predict that high particle susceptibilities of 700 (/1500) are possible for randomly oriented (/uniaxially aligned) 201 nm diameter FeNi3 particles, together with high-frequency stability, shown by low out-of-phase component at 2 MHz. Our findings imply that materials containing nanoparticles have the potential to be tuned to outperform state-of-the-art ferrite inductor core materials at MHz-frequencies.
@@ I Introduction
Magnetic components, such as inductors and transformers, are essential for power electronics in many handheld devices. However, realisation of efficient micro-inductors is challenging, as inductance scales with size [1, 2]. Although a decrease in inductance can be counteracted by higher operating frequencies, magnetic materials available today become inefficient and heat up rapidly at elevated frequencies due to eddy-current losses [3, 4, 5]. Miniaturisation of magnetic components in electronics is therefore limited by the performance of soft magnetic materials [6]. Several roadmaps for power electronics identify the lack of suitable magnetic materials as one of the major obstacles on the way to achieving smaller, faster, greener, and more efficient electronic devices [1, 6].
@@
The challenge lies in achieving sufficiently high magnetic susceptibility while avoiding significant power losses from eddy currents and magnetic hysteresis at high operation frequencies. To this end, for a magnetic material to be used as a core for micro-inductors, its in-phase susceptibility needs to be above 50-100 (the shaded region shown in Fig. 1), while losses should remain below 200 mW/cm3 at an operating frequency of 2 MHz and an induced flux density of 30 mT [3]. If losses are reduced below 20 mW/cm3, a -range of 20-50 would be acceptable [1, 6].
@@
For today’s power converters operating at up to about 0.5 MHz the typical core materials are sintered, fine-grained ceramics of ferrites (MnZnNiFe-oxides), with poor conductivity that limits eddy current losses. The properties of the ferrites vary with their grain sizes and material composition. For example, the commonly used ferrite TDK PC200 has a bulk susceptibility of 800 (shown in Fig. 1) but only up 1-2 MHz, where eddy current losses eventually increases drastically and the susceptibility drops [3]. New ferrite materials, like Fair-Rite 67 (also shown in Fig. 1), enable higher operation frequency, but have lower susceptibilities [4]. So far, attempts to make inductor core materials to cover applications in the high-frequency high-susceptibility regime (the shaded area of interest in Fig. 1) have not been successful [5] and therefore the field of power electronics lacks alternative material solutions.
@@
Magnetic composites using nanoparticles have recently emerged as research materials [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], but their potential as inductor materials has not yet been fully explored. The two main attractive reasons for using magnetic nanoparticles to obtain high susceptibility and low power losses are: 1) the nanoparticles, even if metallic, are so small that eddy currents are negligible, and they can be electrically insulated by a non-conductive matrix, 2) the potential high susceptibility and absence of hysteresis within their superparamagnetic regime could be exploited. Superparamagnetism arises when the magnetic moments of the nanoparticles fluctuate on a faster timescale than the applied field. This depends on particle size, shape, anisotropy, temperature, applied field frequency, and also magnetic interparticle interactions [19].
@@
Several groups have reported susceptibility data for materials containing magnetic nanoparticles ( nm in diameter), as presented in Fig. 1. The measured susceptibilities are generally too low to match the area of interest for future power electronics (the shaded area in Fig. 1). However, it has not been clarified whether composites containing superparamagnetic particles could reach the desired susceptibility in the MHz-range if they were optimized for it. Recently, it has been shown that the susceptibility of single-domain nanoparticles is not limited by their shape (in contrast to soft multi-domain micro-particles) [20]. Hence, there is in principle no upper limit for the effective susceptibility of such nanoparticles, but its actual value will depend on material parameters such as magnetic anisotropy and saturation magnetization as well as particle size and shape.
@@
To evaluate the potential of nanoparticles for use as inductor core materials, we developed a framework for calculating the DC- and AC-susceptibility of single-domain magnetic nanoparticles, both superparamagnetic and blocked. We derive the susceptibility using reported parameter values for specific materials, and considering both uniaxial and cubic anisotropy, combined with the size, shape, and alignment of the nanoparticles. Our theoretical framework reproduces the nanocomposite results shown in Fig. 1, and based on this predictive power, it provides a way forward to tune and optimize properties and performance to reach the region of interest (Fig. 1).
@@ II Theoretical Framework
We consider a stationary single-domain ferromagnetic particle embedded in a solid, non-magnetic matrix. For single-domain particles, atom spins align ferro/ferrimagnetically and rotate coherently such that the magnetic moment of the particle is , where is the particle volume and its saturation magnetization.
@@ II.1 Energy considerations
We define here the various energy terms used later for susceptibility derivations.
@@
For cubic anisotropy we orient the coordinate system based on the case where the cubic anisotropy constant is positive, K 0, i.e., 3 mutually orthogonal easy axes, see Fig. 2(b) but consider both K 0 and K 0 in the following derivations. One easy axis is set to the polar axis and the other two easy axis are along the direction where the azimuth is 0 and . The anisotropy energy can then be expressed as
@@
This takes the same form as (2) for the uniaxial case when the vector scalar product is performed. The distinction of the angles are, however, of importance for the angular integrations in (8) and therefore the vector components are written out here.
@@
with an effective shape anisotropy constant, . and are the demagnetisation factors along the principal spheroid axes. is positive (/negative) for prolate (/oblate) spheroids, respectively, and is the angle between the magnetic moment and the longer (/shorter) principal spheroid axis. For most soft magnetic materials, shape anisotropy dominates over magneto-crystalline anisotropy if the length difference between the axes is larger than 5-10%.
@@
Here is the center-center distance of the two spheres, denotes its Cartesian components, and the subscript for the angles denote the particle number.
@@ II.2 Susceptibility calculations
We derive the susceptibility of single domain particles for the two limiting cases: blocked particles with no thermal agitation, and superparamagnetic particles with their magnetic moments in rapid thermal fluctuation. The characteristic timescale, , for relaxation of the moment between easy directions, depends exponentially on the anisotropy barrier over the thermal energy . Magnetic particles can thus be classified depending on the timescale of superparamagnetic relaxation versus the experimental timescale.
For blocked particles, thermal fluctuations are slower than the experimental timescale. In case of uniaxial anisotropy, the susceptibility depends on the direction of the applied field with respect to the particle easy axis, as described by the Stoner-Wohlfarth model [21]. For randomly oriented particles, the orientation averaged particle susceptibility is
since and and the blocked susceptibilities are and [22]. Blocked particles with uniaxial anisotropy axis aligned parallel to the applied field feature a square hysteresis loop with coercive field of H. If the easy axis is perpendicular to the field, instead, no hysteresis is observed. For blocked particles with cubic anisotropy, a hysteresis loop opening is always observed no matter the direction of the applied field with respect to the 3 or 4 easy axes and thus only a small low field susceptibility is expected [23, 24, 25].
@@
where the integral over indicates integration over all possible energy states . The relevant single particle energies presented in Eqs. (1)-(5) depend on moment direction, thus the integral in Eq. (8) should be carried out over all possible moment directions: . To ease notation we define following energy ratios
@@
The mean magnetisation along the direction of the applied field is then . The susceptibility is found as the derivative of the mean magnetisation with respect to the applied field amplitude
For the non-interacting, uniaxial anisotropy particle case, with the applied field at an angle to the easy axis, one finds the superparamagnetic particle susceptibility to be
@@
Figure 3 shows a plot of Eq. (12) for different values of . The susceptibility ranges from to 0 in case of large anisotropy (). For low anisotropy () goes towards the random case value for all . Special cases of (12) are found in [26, 27, 28, 29, 30, 31].
The orientation average of eq. (12) is the well known superparamagnetic susceptibility
For a non-interacting particle with cubic anisotropy we find no dependence of the susceptibility on easy axis direction with respect to the applied field (i.e. on the angles and ). The initial susceptibility for the cubic anisotropy case is always the same as for the randomly oriented uniaxial anisotropy case in the superparamagnetic regime (Eq. (14)), no matter the strength of the anisotropy constant. This can be explained by the symmetric distribution of energy minima along the 3 or 4 easy axes, which does not change the probability of the moment to point along any specific direction, contrary to the uniaxial case where there is 1 direction (or plane) that is statistically more likely.
Dipolar interactions can be included in the above derivations. For the two particle case, assuming low anisotropy particles (), we find that the susceptibility per particle in case of interactions, , when the two particles form a dimer oriented parallel or perpendicular to the field direction, is, respectively
@@
The dependence of the susceptibility on applied field frequency and the crossover from superparamagnetic to blocked regime can be described and analysed via the concept of AC-susceptibility [22, 26]. At small , where the anisotropy barrier is not affected significantly by the applied field, the AC-susceptibility for a sinusoidal field with angular frequency , , is connected to the superparamagnetic particle susceptibility and the superparamagnetic relaxation time as in the Debye model
@@
For uniaxial anisotropy particles depends on the direction of the applied field with respect to the easy axis [27, 28, 22, 26]. The magnetic moment relaxation time perpendicular to the anisotropy axis is assumed to be short, on the timescale of the attempt time s [19]. The relaxation time over the anisotropy barrier , (parallel to the anisotropy axis), is slower and often described by Arrhenius-type expressions. The AC-susceptibility for randomly oriented uniaxial anisotropy particles, , can thus be expressed by the weighted sum of the AC-susceptibility parallel and perpendicular with the easy axis:
@@
The out-of-phase component gives us an estimate of the hysteresis loss per particle volume by for small applied field amplitudes . For low anisotropy particles the out-of-phase component of the AC-susceptibility is overestimated by the treatment presented above, and should be regarded as an upper bound [32].
@@
Expressions similar to the parallel case of (18) exist for cubic anisotropy, but with a lower effective anisotropy constant of for and for [29, 33, 30, 34].
@@
The formalism introduced above enables the estimate of the maximum susceptibilities of nanoparticles, which is the topic of the next section. We consider candidate materials to be used with applied field at frequencies of up to 10 MHz.
III Particle optimization for nanocomposites
@@
Figure 4 shows the effective in-phase susceptibility of magnetic nanoparticles as function of size, from the superparamagnetic regime to the blocked regime, for randomly oriented, non-interacting, spherical nanoparticles. We have considered a field frequency of 10 MHz and the particle susceptibilities are calculated from Eqs. (12) and (17)-(18). Particle materials used are maghemite (-Fe2O3), Fe, Ni, FeNi3, and FeCo. Despite several of these materials posses cubic crystallinity, effective uniaxial anisotropy and saturation magnetization values from the literature for nanoparticle materials have been adopted [35, 36, 37, 38, 39]. The assumed values for and are given in the figure caption. We have used values reported for 3-10 nm sized particles (21-50 nm for FeNi3) and not considered the possible dependence of saturation magnetisation and effective anisotropy constant on synthesis routine and particle size.
@@
Figure 4 shows that the particle susceptibility increases with particle size for small particles in the superparamagnetic regime. When the particles reach sizes where their superparamagnetic relaxation time is comparable to the timescale of the applied field, the susceptibility peaks. For larger particles, the susceptibility declines towards the blocked value, which depends on particle saturation magnetisation and anisotropy, but not on the particle size, (see Eq. (7)). Fig. 4 illustrates how both FeCo and FeNi3 particles feature effective particle susceptibilities above 100 in the superparamagnetic region.
@@
Our results indicate that nanoparticles with high saturation magnetization and low anisotropy have the largest susceptibility, as their magnetic moments are larger and their transition to the blocked regime occurs at larger diameters. The superparamagnetic region towards the susceptibility peak is of particular interest for applications due to the potential combination of high-susceptibility and limited hysteresis loss. Hence, the optimal size for a given material is on the ascending part of the susceptibility curve, before reaching the maximum of the susceptibility. Based on Fig. 4, we suggest the use of 9 nm FeCo particles or 20 nm FeNi3 particles for high-susceptibility applications such as micro-inductor core materials.
@@
The susceptibility of around 130 for the blocked FeNi3 particles, as seen in Fig. 4, could in principle be of interest, as it may be sufficient for use in micro-inductors. However, blocked particles require precise alignment as only the case shows no hysteresis losses.
@@
Eqs. (17)-(18) reveal that also the dynamic susceptibility depends on the direction of the particle with respect to the applied field. Fig. 5 shows the frequency dependence of the in- and out-of-phase susceptibility for uniaxial anisotropy FeNi3 particles with log-normal distributed diameters of 201 nm for cases with anisotropy axes parallel, perpendicularly and randomly oriented with respect to the applied field. In the frequency range of 10 kHz to 10 MHz the 20 nm FeNi particles aligned with their easy axis parallel to the applied field have an in-phase susceptibility of 1600, i.e., more than twice the value for the randomly oriented case (700) and more than six times the perpendicular ”hard-axis” case (230). At around 100 MHz the in-phase part drops and the out-of-phase component peaks for the parallel aligned and random cases. This is due to in this frequency range. For the random case the susceptibility drops towards the random orientation blocked susceptibility of Eq. (7). For the perpendicular case we observe a completely flat susceptibility up to the GHz regime where , and we find that the susceptibility does not depend on particle size, with values close to the blocked, aligned case of Eq. (7). Alignment of particles by easy/hard axis thus allows for tuning of the magnetic properties.
@@
Even though shape anisotropy can increase susceptibility for particles with their easy axis aligned to the applied field, (see Eq. (12) and Fig. 3), the anisotropy enhancement will increase relaxation time for such particles and thus the out-of-phase component, corresponding to power losses due to magnetic hysteresis. This is illustrated in Fig. 6, where the normalised in- and out-of-phase components of the susceptibility are shown for 201 nm FeNi3 particles aligned with easy-axes along the applied field. We have assumed that shape and magneto-crystalline anisotropy are in the same direction, i.e. . Fig. 6 shows that the in-phase component of the susceptibility is relatively insensitive to a shape anisotropy increase (0.3 % at ), while the increase in out-of-phase component is substantial (22 % at ). We also observe that for larger shape anisotropy () the particle becomes blocked. Particles smaller than the blocking diameter at a given frequency will show a more significant increase in in-phase susceptibility. However, the connected increase of the out-of-phase susceptibility will be similar to the case in Fig. 6. These results on FeNi3 particles lead us to conclude that spherical particles are the most optimal, as any shape anisotropy present increases the out-of-phase component of the susceptibility due to increased superparamagnetic relaxation time, resulting in larger losses per induced magnetisation as discussed in the next section.
@@
Hysteresis losses per particle can be calculated from the out-of-phase component of the susceptibility [Eqs. (12),(17) and (18)]. For nanocomposites containing superparamagnetic particles we assume a linear dependence of nanocomposite susceptibility on the volume fraction and particle susceptibility. Therefore the applied field strength needed to obtain a certain magnetic flux in the material varies between materials. Hence, for a given application with a required flux density, the loss might be larger for a low susceptibility material than for high susceptibility materials due to the need for larger applied field. Therefore, assuming no interaction and same particle characteristics, the losses will increase with decreasing volume fraction of particles when the nanocomposites generate equal flux density.